OpenStudy (anonymous):

Find the point on the terminal side of θ = pi/4 that has an x coordinate of 2. Show your work for full credit.

OpenStudy (anonymous):

when x=2, y=? tan(theta)= y/x solve for y by plugging in theta as pi/4 and x as 2 right?

OpenStudy (anonymous):

tan(pi/4) = y/2 tan(pi/4) =1 1 = y/2 *2 y =2

OpenStudy (anonymous):

thank you so much